Calculate V-3+ 3i. Give your answer in a + bi form. In polar form, use the angle 0 <o<27. Submit Question
Answer #1

Solutions we evaluate folos form of 5.331 in form of We ving atli Where the angle o LOL 20 80-an.com MEZ and -3 cio 352 = groOos arcos(-a) - and m-a 315) 3J-3439 = (arba) 3 81 4+0) - ) 31-343} = (353) (cos Po ti sinf 43 R = (312) Radius atib Corel ta